x/2 + 2/3=3/4 I am new to this and am stuck on this question. Thank you for your help!! Belle
Answer by fractalier(6550)

The easiest way to handle fractional equations is to multiply everything by the lowest common denominator first...here that would be 12...so we have
x/2 + 2/3 = 3/4
12[x/2 + 2/3 = 3/4]
6x + 8 = 9
now subtract 8 from both sides and then divide by 6 and get
6x = 1
x = 1/6